Cellular T-equivariant cohomology

Matija Cencelj, Ljubljana, Slovenia

Keywords: equivariant cohomology, G-complexes, cellular cohomology AMS Subject classification: 55N91

Abstract: Let X be a (relative) T-complex with a compatibile CW filtration, T being the circle. We show that Borel T-equivariant cohomology with coefficients in Z (or a field of characteristic 0) of X is the homology of a double complex of cellular cochains.